# SonifyAR: Context-Aware Sound Generation in Augmented Reality

### v1

@@ 6. Applications
To demonstrate the potential and technical capability of SonifyAR, we explore five example application scenarios that utilize SonifyAR’s sound authoring pipeline in AR applications. All examples are also shown in supplementary videos for better presentation.
@@ 6.1. Education
AR have been widely explored in STEM education (sirakaya2022augmented, ). As existing educational applications try to simulate and visualize physical or chemical phenomenon in AR (radu2019can, ; irwansyah2018augmented, ), we envision that SonifyAR can help add sound into the simulation and improve the blending between real environment and virtual contents.
To showcase such possibility, we implement an AR physics experiment with SonifyAR demo application. We create a 3D model to show the conservation of momentum, which is a slope and two metal balls located at top and bottom of the slope (Figure 8). Upon initiating the experiment, the top ball rolls down the slope and collides with the bottom ball. This collision causes the bottom ball to move forward and eventually fall, while the top ball comes to a stop. With SonifyAR, collision events are automatically detected, and material-specific sound effects, such as the clashing of two metal balls or a metal ball dropping onto table, are seamlessly generated. The resulting sound effects can be experienced in the supplementary video.
@@ 6.2. Accessibility
AAR (Audio Augmented Reality) has been explored as accessibility assistance for blind or low vision people (ribeiro2012auditory, ; coughlan2017ar4vi, ). By sonifying real world environments and virtual contents, blind or low vision (BLV) people can better interpret visual information in both reality and virtuality. Following this thread of work, we envision SonifyAR being able to assist on accessibility in AR experiences in three ways. First, by supporting AR sound authoring, SonifyAR encourages creators to add sound effects in AR experiences, which can help BLV users consume visual contents. Second, by providing context-aware AR sound in 3D audio, people with low-vision can better navigate virtual objects in AR environment. Third, by enabling user to interact with real-world surfaces via AR (e.g. tapping on a real world surface), user can explore the surrounding space via the AR interface.
@@
We use a toy robot 3D model and the SonifyAR demo app to showcase the accessibility benefits. The toy robot (Figure 9) can be placed on any indoor surface and has a walking animation. By authoring different bouncing sound effects for different surfaces (e.g. wood and carpet), user can better perceive the location of the robot while it is walking.
@@ 6.3. Sonify Existing Apps
Due to the extensibility of the text template used in SonifyAR’s sound acquisition pipeline, we foresee its application in existing AR apps if integrated as an extension. If implemented at the SDK level (e.g. ARKit and ARCore), SonifyAR can automatically capture textual description of AR events and provide sonification using the automatic sound acquisition results.
@@
We explore this application possibility with a Wizard of Oz (WoZ) prototype on a phone-based AR application called ARVid (arvid, ), which is an AR app that puts 3D animated objects into real world environment. We select a dinosaur animation as an example (Figure 10), and conceptualize that SonifyAR pipeline is adopted into the AR infrastructure. By manually feeding the context information (e.g. “The dinosaur is eating”) into the SonifyAR pipeline, we generated sound assets that well-match the 3D animations. The results are shown in the supplementary video.
6.4. Sonify virtuality-reality interaction on AR/MR headset
As recent AR/MR headsets become increasingly popular in entertainment and productivity tasks, manufacturers have established safety guidelines, such as setting up safety zone and clear up indoor spaces. We envision SonifyAR contributing to this topic by creating material-aware sound effects during headset usage when application windows or virtual object intersect with real world surfaces or objects.
@@ 6.4. Sonify virtuality-reality interaction on AR/MR headset
We implement a WoZ prototype based on Apple Vision Pro’s user interaction (Figure 11). When using Vision Pro, user can actively adjust the position of application windows by hand gestures. SonifyAR can enhance this interaction by generating AR sounds when virtual windows are placed on or intersect with indoor surfaces like walls, desks, and floors. By processing these interactions and applying sound transfers with text prompts such as “a virtual window bumps into a wall” and “a virtual window collides with a carpet”, SonifyAR aims to improve the permanence of virtual windows and increase awareness of surrounding objects as users navigate through virtual windows. See supplementary video for more direct inspection of the sound effects.
6.5. Sound asset generation for traditional AR authoring process
Besides serving as a PbD authoring experience, SonifyAR’s sound authoring pipeline also have the potential of improving existing PbS AR authoring tools in sound assets provision. As the mainstream AR authoring tools like Adobe Aero and Apple Reality Composer already represent their AR interaction as text like “Tap & Play Sound” and “Proximity & Jiggle”, these specifications can be fed into SonifyAR’s text-based sound authoring pipeline. In this case, SonifyAR works as a plugin of these tools and provide matching sound assets for AR interactions.
@@ 6.5. Sound asset generation for traditional AR authoring process
We implement a WoZ prototype using Apple’s Reality Composer and a 3D model of an animated machine (Figure 12). When authoring the “Play Sound” behavior of this model, context information like the model description, triggering condition, and animation description, can be fed into SonifyAR’s text-based authoring pipeline and lead to two sound assets from generation and retrieval methods. These two sound assets will appear at top of the sound asset selection as “SonifyAR Recommended”. See supplementary video for more details.
@@ 7. Discussion and Future Work
In this paper, we designed and implemented an AR sound authoring pipeline that automatically generates sound assets based on context information. We evaluated and discovered the opportunities in the AR sound interaction space, and implemented an LLM+Generative AI pipeline for sound acquisition. We tested the pipeline with a usability study on our prototype and further discussed the application opportunity with five examples.
@@ 7.1. Limitations and failure cases
While SonifyAR was positively evaluated in our user study and demonstrated broad application potential, we observed limitations in the system adaptability and encountered failure cases in multiple aspects.
Firstly, since the sound acquisition pipeline relies on LLM as the context collector and overall controller, errors of the LLM can lead to failure cases in sound acquisition. For example, the LLM can extract the wrong keywords in the context information and prompt sound generation model with inaccruate text input (e.g. LLM provide prompt of “sliding plastic on carpet” while the sliding model is actually a metal toy robot) that lead to non-matching outputs.
Secondly, although AudioLDM (liu2023audioldm, ) is the state-of-the-art text2audio generator and performs decently in most cases, it can sometimes produce subpar outputs that are noisy or do not match prompts. Based on our experiences, such inconsistency in performance can be hard to predict thus cannot be completely solved by prompt engineering.
Thirdly, some AR interactions can be highly time-sensitive (e.g. objects colliding, the playing of animations that shows very clear action sequences) and require sound assets being closely synchronized with visual contents. Currently, we provide example sound assets for AR interactions of “Colliding”, “Tapping” and “’Sliding’. By conducting sound style transfer using well-timed example sounds as inputs, the output sounds can be formatted to be in sync with the AR interaction. However, for animations, we currently do not have methods to sync the generated sound with the key frames in animation, so some generated sounds can be off-sync with animation. We believe this 3D animation-to-audio synthesis challenge can be a standalone topic for future researchers to solve.
7.2. Expansion of framework
Currently, we implement the automated AR sound acquisition with six example AR event types, as listed in section 4. This selection is primarily for proof of concept. Future work, or other systems adopting SonifyAR’s text-based sound acquisition pipeline, could easily incorporate additional types of AR interactions, provided they have clear and concise text descriptions. For example, event types used in Adobe Aero and Reality Composer, such as “Spin”, “Emphasize”, “Expand”, or “Jump”, can be well-incorporated into the text template of SonifyAR. The flexibility and extensibility of our text-based pipeline is also explored in our application examples.
@@ 7.3. Future work
First, our current SonifyAR app is a simplified tool that exclusively supports the authoring of AR sound and uses pre-crafted AR scenes as input. For wider applicability that reflects the real-world needs of AR content creation, we recognize the need to incorporate the AR sound authoring pipeline into the entire AR content creation process. Our future goals include integrating the SonifyAR framework with established tools like Reality Composer and Adobe Aero.
Secondly, SonifyAR adopts a straightforward process of using a text template to compile context information collected from multiple sources, including virtual object semantics and real world semantics. Currently, we use one material segmentation model to perceive real-world context information. In future developments, we aspire to incorporate more sophisticated computer vision models, or multi-modal foundational models, for deeper understanding of both virtual objects and real world environment, enhancing the context acquisition process with greater precision and broader adaptability.
Lastly, the current sound generation output of SonifyAR presents opportunities for further refinement. Recognizing the rapid advancements in the AI research domain, our system has been designed as a modular framework, ensuring that as more advanced models emerge, they can be seamlessly integrated. Every model (e.g., LLM, CV, Sound Generation) within SonifyAR is replaceable, allowing SonifyAR to easily adapt and stay at the forefront of innovation in this field.
@@ 8. Conclusion
In this paper, we present SonifyAR, a context-aware sound authoring system designed for sonifying AR events. Our work implements a PbD authoring pipeline which enables sound asset acquisition using context information. With SonifyAR, users could demonstrate AR interactions and have AR sound assets automatically acquired. Our studies validate the usability and overall performance of our system. Our five application examples further widens our application potential. We believe SonifyAR can be applied in application fields like education, BLV assistance, and AR headset safety. We also envision it incorporated into existing AR authoring tools to strengthen their sound authoring capabilities. This research advances literature in AR sound authoring, while also opening up new research avenues for the application of LLM and generative models in AR systems.

### v3

@@ 6. Applications
To further explore and demonstrate the potential of our approach, we authored five AR scenarios with SonifyAR. See the supplementary video for additional details and an audio-visual demonstration of the resulting SonifyAR sound effects.
@@ 6.1. Education
AR has been widely explored in STEM education (e.g., (Sırakaya and Alsancak Sırakaya, 2022; Kang et al., 2020, 2019, 2016; Suzuki et al., 2020; Chulpongsatorn et al., 2023)). As existing educational applications try to simulate and visualize physical or chemical phenomenon in AR (Radu and Schneider, 2019; Irwansyah et al., 2018), SonifyAR can improve the blending between the real environment and virtual content through immersive, appropriate sound effects. To showcase such a possibility, we implemented an AR physics experiment using SonifyAR. To help illustrate one of Newton’s laws of motion–the conservation of momentum—we created an AR scene with a downward ramp and two metal balls: one at the top and the other at the bottom (Figure 8). Upon initiating the experiment, the top ball rolls down the slope and collides with the bottom ball. This collision causes the bottom ball to move forward and eventually fall, while the top ball comes to a stop. With SonifyAR, collision events are automatically detected and material-specific sound effects, such as the clashing of two metal balls or a metal ball dropping onto table, are seamlessly generated.
@@
[…]
A figure showing an AR physics experiment and SonifyAR’s authoring results. A: a virtual slope with two metal balls on it. The top ball slide down and collide with the lower one. Then the lower one drop down to the table while the top one remain on slope. B and C: the context information of the two collide events; D: the resulted sound assets shown in UI.
@@ 6.2. Accessibility
AAR (Audio Augmented Reality) has been explored as accessibility assistance for blind or low vision people (Ribeiro et al., 2012a; Coughlan and Miele, 2017). By sonifying real world environments and virtual contents, blind or low vision (BLV) people can better interpret visual information in both reality and virtuality. We envision SonifyAR assisting AR accessibility in three ways. First, by supporting AR sound authoring, SonifyAR encourages creators to add sound effects in AR experiences, which can help BLV users interpret visual content. Second, by providing context-aware AR sound in 3D audio, people with low-vision can better navigate virtual objects in the AR environment (Ribeiro et al., 2012a). Third, by enabling user to interact with real-world surfaces via AR (e.g. tapping on a real world surface), user can explore the surrounding space via the AR interface.
@@
[…]
A figure showing an AR robot walking and SonifyAR’s authoring results. A: a virtual robot walk on wooden table. B: the context information of the walking; C: a virtual robot walk on carpet surface. D: the context information of the walking; E and F: the resulted sound assets shown in UI.
To showcase potential accessibility benefits, we use a toy robot 3D model (Figure 9). As the robot virtually walks from one physical surface to another (e.g., from a wood floor to a carpeted floor), the auto-generated sound effects change appropriately. In this way, a BLV user can better perceive the location and activity of the robot as it is walking.
@@ 6.3. Sonify Existing Apps
Due to the extensibility of the text template used in SonifyAR’s sound acquisition pipeline, it can be easily implemented into existing AR apps. For example, if implemented at the SDK level (e.g. ARKit and ARCore), SonifyAR can automatically capture textual description of AR events and provide sonification using the automatic sound acquisition results.
@@
[…]
A figure showing an dinosaur roaring and SonifyAR’s authoring results. A: a virtual dinosaur roar in an AR scene; B: the context information of the roaring; C: the resulted sound asset.
We explore this application possibility with a Wizard of Oz (WoZ) prototype on a phone-based AR application called ARVid (arv, [n. d.]), which is a downloadable AR app that puts 3D animated objects into real world environment. We selected a dinosaur animation as an example (Figure 10) and use WoZ to simulate the direct integration of the SonifyAR pipeline into ARVid app. By manually feeding the context information (e.g. “The dinosaur is eating”) into the SonifyAR pipeline, we generated sound assets that well-match the 3D animations.
6.4. Using SonifyAR on an MR Headset
As recent AR/MR headsets become increasingly popular, manufacturers have established safety guidelines, such as setting up safety zone and clear up indoor spaces. We envision SonifyAR contributing to this topic by hinting users of the existence of real-world entities with material-aware sound effects when application windows or virtual object intersect with real world surfaces or objects.
@@ 6.4. Using SonifyAR on an MR Headset
[…]
A figure showing Vision Pro’s interface and SonifyAR’s authoring results. A: a virtual browser window rests on a sofa in Vision Pro’s interface; B: the context information of the collision; C: the resulted sound asset.
We implement a WoZ prototype based on Apple Vision Pro’s user interaction (Figure 11). When using the Vision Pro, a user can actively adjust the position of application windows by hand gestures. SonifyAR can enhance this interaction by generating AR sounds when virtual windows are placed on or intersect with indoor surfaces like walls, desks, and floors. By processing these interactions and applying sound transfers with text prompts such as “a virtual window bumps into a wall” and “a virtual window collides with a carpet”, SonifyAR aims to improve the permanence of virtual windows and increase awareness of surrounding objects as users navigate through virtual windows.
6.5. Augmenting AR Authoring Processes
Besides serving as a standalone PbD authoring experience, SonifyAR could improve existing AR authoring tools. As mainstream AR authoring tools like Adobe Aero and Apple Reality Composer already represent their AR interaction as text like “Tap & Play Sound” and “Proximity & Jiggle”, these specifications can be fed into SonifyAR’s text-based sound authoring pipeline. In this case, SonifyAR works as a plugin and provides matching sound assets for AR interactions.
@@ 6.5. Augmenting AR Authoring Processes
[…]
A figure showing Reality Composer’s interface and SonifyAR’s authoring results. A: Reality Composer’s mobile interface with a sewing machine in the canvas; B: Reality Composer’s UI for authoring the ”play sound” action of the virtual object; C: the context information of the ”play sound” action; D: SonifyAR’s generation result as a sound asset; E: SonifyAR’s retrieval result as a sound asset; F: the WoZ interface of Reality Composer showing SonifyAR’s sound assets.
We implemented an additional WoZ prototype using Apple’s Reality Composer and a 3D model of an animated machine (Figure 12). When authoring the “Play Sound” behavior of this model, context information like the model description, triggering condition, and animation description, can be fed into SonifyAR’s text-based authoring pipeline. The resulting retrieved/generated sound assets appear at the top of the authoring UI as “SonifyAR Recommended”. See supplementary video for more details.
@@ 7. Discussion and Future Work
In this paper, we designed and implemented an AR sound authoring pipeline that automatically generates sound assets based on context information. We evaluated and identified opportunities in the AR sound interaction space, and implemented a custom LLM+Generative AI pipeline for sound acquisition. We tested the pipeline with a usability study and through a ”proof-by-demonstration” across five application scenarios.
@@ 7.1. Limitations and Failure Cases
While SonifyAR was positively evaluated in our user study and demonstrated broad application potential, we observed limitations in the system adaptability and encountered some failure cases.
Firstly, since the sound acquisition pipeline relies on an LLM as both the context collector and overall controller, LLM errors can lead to failures. For example, the LLM can hallucinate about the context information and prompt sound generation model with inaccruate text input (e.g., LLM provides prompt “sliding plastic on carpet” when the provided context information is the virtual model, a metal toy robot, slides on wooden floor) that lead to non-matching outputs.
Secondly, although AudioLDM (Liu et al., 2023) is a state-of-the-art text-to-audio generator and performs well in most cases, it can sometimes produce subpar outputs that are noisy or do not match prompts. Based on our experiences, such performance inconsistency can be hard to predict thus cannot be completely solved by prompt engineering.
Lastly, some AR interactions can be highly time-sensitive (e.g., objects colliding, playing animations that show clear action sequences) and require sound assets to be precisely synchronized with visual content. Currently, we use two methods to synchronize sounds with AR events. First, we provide example sound assets for AR interactions of “Colliding”, “Tapping” and “’Sliding’. By conducting sound style transfer using well-timed example sounds as inputs, the output sounds can be formatted to be in sync with the AR interaction. Secondly, we bind colliders to 3D model skeleton joints, ensuring that the generated sound is played only when a collision is detected between an animated model and real-world surfaces. However, these synchronization method do not apply to 3D assets without skeletons or sounds unrelated to collisions, such as dinosaur roaring and mechanical arm unfolding. Fully addressing this 3D animation-to-audio synchronization challenge is beyond our current scope.
@@ 7.2. Future Work
First, our current SonifyAR app is a simplified tool that exclusively supports the authoring of AR sound and uses pre-crafted AR assets as input. For wider applicability that reflects the real-world needs of AR content creation, we recognize the need to incorporate the AR sound authoring pipeline into the entire AR content creation process. Our future goals include integrating the SonifyAR framework with established tools like Reality Composer and Adobe Aero, and also building a Unity SDK to enable usage of our system among professional AR developers. We will also expand the list of AR event types supported by SonifyAR, adding categories such as “Spin”, “Emphasize”, “Expand”, or “Jump”—all of which can easily be incorporated into the text template of SonifyAR.
Secondly, SonifyAR adopts a straightforward process of using a text template to compile context information collected from multiple sources, including virtual object semantics and real world semantics. Currently, we use one material segmentation model to perceive real-world context information. In future developments, we aspire to incorporate more sophisticated computer vision models or multi-modal foundational models, for deeper understanding of both virtual objects and real world environment, enhancing the context acquisition process with greater precision and broader adaptability. Thirdly, SonifyAR has some basic error-handling features, such as the automatic removal of failed retrieval results and editing options for subpar sound assets (Section 4.3). Additionally, users can correct recognition errors, such as incorrectly recognized materials, by specifying the correct material with text. In future versions of SonifyAR, we plan to incorporate more advanced error-handling mechanisms, such as deploying a sound quality inference model to validate the output and automatically re-prompt the sound generation model when subpar sounds are generated.
Lastly, the current sound generation output of SonifyAR presents opportunities for further refinement. Recognizing the rapid advancements in AI, our system has been designed as a modular framework, ensuring that as more advanced models emerge, they can be seamlessly integrated. Every model (e.g., LLM, CV, Sound Generation) within SonifyAR is replaceable, allowing SonifyAR to easily adapt and stay at the forefront of innovation in this field.
@@ 8. Conclusion
In this paper, we present SonifyAR, a context-aware sound authoring system designed for sonifying AR events. Our work implements a custom PbD authoring pipeline which enables sound asset acquisition using context information and AI. With SonifyAR, users demonstrate AR interactions and have AR sound assets automatically acquired. Our studies validate the usability and overall performance of our system. Our five application examples further supports the potential of our approach. This research advances literature in AR sound authoring, while also opening up new research avenues for the application of LLM and generative models in AR systems.
Acknowledgements.
This research was mainly conducted during an internship at Adobe Research and also supported by NSF award #1763199. We thank Dr. Cuong Nguyen for his suggestions to our research and our user study participants for their participation.
Appendix A Task Prompts
Here we provide text prompts for tuning a GPT4.0 model into our sound acquisition process controller.
System Context: You are an assistant helping user generate or retrieve matching sound assets for augmented reality events. You will be provided with the description of the event, including: who caused this event, what is this event, and what is the subject of this event, which can be null for some cases. Note that the description of the causer and subject of event can sometimes be long and you need to extract the important part of it to create concise replies. For example, when being told that the causing entity of the event is “metalball of a plastic slope with balls running on it” you should know that the only mattering keyword here is the metalball. You have three methods to provide sound assets. Method 1 is recommending from a predefined list of sound assets, each with a name describing the content. You will recommend the best matching name based on event description. The full list of sound asset will be provided at the end of this context; Method 2 is an online sound retrieval API, for which you should generate a simplified text prompt to search with, preferably within 4 words; Method 3 is a diffusion model that takes text prompt and generate sound assets. This method has three functions: generating new sound file with text prompt, generating similar sound file based on an existing sound, and transferring an existing sound with a text prompt. For each conversation, you will first provide a response about all three sound sources. User may ask following up questions in reply, then you just reply based on specific feedback. When providing results, you should always follow a strict format starting and end with #. For method 1, you should reply the best matching sound asset in #method1:FILENAME# format. For method 2, you should reply in format of #method2:PROMPT#, please replace the PROMPT with your generated prompt. For method 3, when generating new sound, you should reply with #method3generation:PROMPT#, please replace the PROMPT with your generated prompt; when generating similar sound of a existing sound, reply with #method3similar#, when transferring an existing sound, reply with #method3transfer:PROMPT#, please replace the PROMPT with your generated prompt. Note that the hashtags need to be kept at both start and end of the result, the FILENAME need to be replaced by the exact name provided in the asset list, with no exception. The PROMPT should be replaced with the prompt you generate for these cases. Don’t explain why you have these prompts, just stricktly follow the formats. Here I list names of sound assets to recommend from, each filename represents its content: 1. Knock surface 2. Sliding sound 3. Crash Bulb Break ¡… 32 assets in total…¿
We also prompt GPT4.0 to better parse the context information of specific AR events. We show one example event in the following prompt:
Conversation Prompt: Please give me the results as hinted by context. For this specific event, the description of the AR event is: The sound-producing AR event is TapSurface. The causing entity of this event is a metal robot. The target entity is a wooden surface. You should reply top5 results with method 1, ensuring that the filename provided is exactly from the asset list. You should also give one result with method 3 transfer, with a reasonable prompt that could transfer an existing tapping sound to a more material-aware sound. For example, for a tapping on a wooden surface, you could use tapping wood as the prompt. Reply with the format of #method3transfer:PROMPT#, replacing the PROMPT with your generated prompt. You can add a ’high quality’ tag in the prompt to help improve the sound quality.